OpenStudy (anonymous):

A 2-mi cab ride costs $5.25. A 5-mi cab ride costs $10.50. How much does a 3.8-mi cab ride cost? A. $14.00 B.$10.75 C.$8.40 D.$6.30

OpenStudy (anonymous):

I would go about solving this by first finding how much it costs fir 1 mile of a cab ride. So divide 5.25 by 2.

OpenStudy (anonymous):

Once you have figured out how much it is for 1 mile, you can multiply that number by how many miles you have. Tell me what you get and how and I will check your work

OpenStudy (anonymous):

Igreen dont jack the post again lol

OpenStudy (anonymous):

I see you lurking

OpenStudy (anonymous):

i tried but if you multiply the cost for one mile by 5 to get the 5 mile then its wrong

OpenStudy (anonymous):

First divide 5.25 by 2 and tell me what you get

OpenStudy (igreen):

JUST SPYING! AHAHAHAHAHAHAHA!

OpenStudy (anonymous):

5.25 divided by 2 is 2.625. but like i said 2.625x5 is 13.125. so if i multiply it by 3.8 ill get 9.975 which is wrong because the 5 mile is wrong.

OpenStudy (anonymous):

Hold up, you went ahead and in the wrong direction. The beginning as good. You found out for 1 mile of a cab ride it costs $2.625

OpenStudy (anonymous):

the 9.975 is correct

OpenStudy (anonymous):

Oh okay this is a completely different problem then I thought, it actually has different costs for different mileages.. Sooo Write a linear equation using the given values as follows x1 = 2; y1 = 5.25 x2 = 5; y2 = 10.50

OpenStudy (anonymous):

but thats none of the options. and how can that be correct if when you multiply the 1 mi by 5 to get the 5 mi its wrong

OpenStudy (anonymous):

Find the slope using slope formula (y2-y1)/(x2-x1)

OpenStudy (anonymous):

Use the point/slope equation: y - y1 = m(x - x1) y - 5.25 = 1.75(x - 2) y - 5.25 = 1.75x - 3.5 y = 1.75x - 3.5 + 5.25 y = 1.75x + 1.75; the equation ; Find the cost of a 3.8 mi cab ride, substitute 3.8 for x and find y y = 1.75(3.8) + 1.75 y = 6.65 + 1.75 y = $8.40 for 3.8 mi ride